OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ dx }{(16+9x^2)^{\frac{ 3 }{ 2 }} }\]

hartnn (hartnn):

when you see the form a^2x^2 +b^2 you plug in x = (b/a) tan u

hartnn (hartnn):

here, put x = (4/3) tan u

OpenStudy (anonymous):

what is the u?

hartnn (hartnn):

new variable after substitution...

hartnn (hartnn):

x = (4/3) tan u dx = ... ?

OpenStudy (anonymous):

4/3 sec^2 du?

hartnn (hartnn):

correct. and what does denominator simplify to ?

OpenStudy (anonymous):

4/3 sec^3du

hartnn (hartnn):

du ? 4/3 ? i just get 64 sec^3 u

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

then where do i go from there?

hartnn (hartnn):

what gets cancelled!?

hartnn (hartnn):

isn't just cos u/48 remain ? do i need to show you the simplification part ?

OpenStudy (anonymous):

well i'm just not thinking this is right because according to the book i should be getting. \[\frac{ x }{ 16\sqrt{16+9x ^{2}} }+C\]

hartnn (hartnn):

we will get to that after simplifying the answer ..... but did u get till cos u/48 ?

OpenStudy (anonymous):

no, i guess not

OpenStudy (anonymous):

where did you go after we got 4/3 scec^2 du.

OpenStudy (anonymous):

you just plugged that in for x?

hartnn (hartnn):

we simplified the denominator....could you ?

hartnn (hartnn):

16 + 9* 16/9 tan^2 u = 16 +16 tan^2 u = 16 sec^2 u that ^2/3 =....

hartnn (hartnn):

16^2/3 = 64 (sec^2) ^ 2/3 = sec^3 now got how the denom = 64 sec^3 x ?

OpenStudy (anonymous):

oh ok I get that

hartnn (hartnn):

so, sec^2 from numerator gets cancelled with sec^3 to get 1/ sec = cos

OpenStudy (anonymous):

Oh i forgot we changed dx

hartnn (hartnn):

till which step did u totally get ?

OpenStudy (anonymous):

so far I've gotten everything once you actually put it in the equation. we should have \[\frac{ (4/3)cosx }{64 }\]

OpenStudy (anonymous):

which simplifies to cos u/48

hartnn (hartnn):

thats cos u/48 so integrate that

hartnn (hartnn):

easiest trigo integral ever :)

OpenStudy (anonymous):

wouldn't that be -sin u/48

hartnn (hartnn):

that would be +sin u/48+c now just put u = tan inverse (3x/4)

OpenStudy (anonymous):

would that be (cos 3x/4)/48

hartnn (hartnn):

no.... (1/48)sin (tan inverse 3x+4) +c that will not be cos

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

\(\Huge \sin (\tan^{-1}A) = \dfrac{A}{\sqrt{1+A^2}}\) here A = 3x/4

OpenStudy (anonymous):

ok. i don't have that on my table of integrals. so i'm going to write that down

hartnn (hartnn):

sure, thats inverse trigonometric relation.

hartnn (hartnn):

(1/48)sin (tan inverse 3x/4) +c = (1/48) [(3x/4) / \sqrt {1+ (3x/4)^2}] +c if you simplify this, you will het the answer same as that given in your book :)